The length of a rectangle is 2 cm more than twice its width. If the perimeter of the rectangle is 52 cm, find the dimensions of the rectangle.
Found 2 solutions by checkley75, rmromero:
Answer by checkley75(3666)

L=2W+2
2L+2W=PERIMETER
2(2W+2)+2W=52
4W+4+2W=52
6W=52-4
6W=48
W=48/6
W=8 ANSWER FOR THE WIDTH. THUS:
L=2*8+2
L=16+2
L=18 ANSWER FOR THE LENGTH
PROOF
2*18+2*8=52
36+16=52
52=52

Answer by rmromero(383)


The length of a rectangle is 2 cm more than twice its width.
If the perimeter of the rectangle is 52 cm, find the dimensions of the rectangle.
What is asked in the problem?
find the dimensions of the rectangle.
Given:
The length of a rectangle is 2cm more than twice its width
The perimeter of the rectangle is 52cm
Representation:
Let x = the width of the rectangle
2x+2 = the length of the rectangle
Equation
2l + 2w = P
2(2x+2) + 2x = 52 Multiply (Distributive Property)
4x + 4 + 2x = 52 Combine like terms
6x + 4 = 52
6x + 4 - 4 = 52 - 4 subtract 4 both sides
6x = 48 divide both sides by 6
x = 8 cm -----> Width
length = 2x + 2
= 2(8) + 2
= 16 + 2
= 18 cm ----- length
Checking
2L + 2W = P
2(18) + 2(8) = 52
36 + 16 = 52
52 = 52